OpenStudy (anonymous):

the weight of 10,000 identical samples of a substance is 1 pound. what is the wight of 100 samples?

OpenStudy (anonymous):

0.2

OpenStudy (anonymous):

It's wrong... Sorry @coasterman777

OpenStudy (whpalmer4):

if 10000 samples = 1 lb, what is the weight of 100 samples? let x be the weight of 100 samples (in lbs): \[\frac{10000 \text{ samples}}{1\text{ lb}} = \frac{100\text{ samples}}{x\text{ lb}}\]Cross multiply to solve for \(x\) \[10000x = 100\]divide both sides by \(10000\)\[x = \frac{100}{10000} = 0.01\]So 100 samples would weigh \(0.01\text{ lb}\).